Astrid1:

y=-3x-6 Hewp me solve

sxdsouls:

its a graph question right ...

bella345:

get x by it's self

Astrid1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sxdsouls its a graph question right ... \(\color{#0cbb34}{\text{End of Quote}}\) Yeah

boredfr:

Your graph should align with this however they might ask you for the slope.

sxdsouls:

breh-

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sxdsouls breh- \(\color{#0cbb34}{\text{End of Quote}}\) :)~~ferny give meh medalll

xxDeppressionxx:

Add 6 to both sides. y+6=-3x Divide both sides by -3 -y+6/3=x Switch sides. x=-y+6/3 ​

supie:

If we're solving for \(x\) The first thing we do is flip the equation so that `y` comes first So \(y=−3x−6\) becomes \(−3x−6=y\) Then we \(+6\) to both of the sides of the equation \(−3x−6+6=y+6=−3x=y+6\) Then, \(÷-3\) to both sides \(\LARGE\frac{−3x}{-3}=\frac{y+6}{-3}\) Then \(\frac{-3x}{-3}\) cancels So all you have left is \(\frac{y+6}{-3}\) which =?

xxDeppressionxx:

i already said something supie

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx i already said something supie \(\color{#0cbb34}{\text{End of Quote}}\) nice